The following chemical reaction is at equilibrium at 25°C in a 1.0 L flask, with [A] = 0.28 M.

A(g) mc028-1.jpg B(g) K = 2.47

Suppose that 0.10 moles of A is added to the system. What will be the concentration of B after equilibrium has re-established
1
Expert's answer
2020-06-04T10:29:24-0400

K = [B] / [A], [B] = K[A] = 2.47 × 0.28 = 0.6916

2.47 = (0.6916 + x) / (0.28 + (0.1 - x))

x = 0.0712

CB = [B] + x = 0.6916 + 0.0712 = 0.7628

Answer: 0.7628 M.